OpenStudy (dumbsearch2):

Sam can paint the house in 8 days by himself, while Betty can paint the same house in 56 days. How long would it take Sam and Betty to paint the house together?

Parth (parthkohli):

If one dude can do a work in \(x\), and the second can do it in \(y\), then both do it in\[\dfrac{1}{x} +{\dfrac{1}{y}}\]

hero (hero):

Let S = the number of days Sam can paint a house alone B = the number of days Betty can paint a house alone Then together, they can paint the house in \(\large\frac{S \times B}{S + B}\) days

hero (hero):

So in essence, compute: \[\frac{8 \times 56}{8 + 56}\]

hero (hero):

@dumbsearch2, did you figure it out?

OpenStudy (dumbsearch2):

112, right? (btw thanks for helping. :))

hero (hero):

No that's not right. Did you compute that by hand or calculator?

OpenStudy (dumbsearch2):

calculator. :)

OpenStudy (dumbsearch2):

and it says 112! :| with what you provided...

hero (hero):

You're inputting it into the calculator wrong.

hero (hero):

Maybe you'll understand what to do if I wrote it this way: (8 × 56) ÷ (8 + 56) Order of operations is important, especially with a calculator.

OpenStudy (dumbsearch2):

youre right! :) the answer is 7?

hero (hero):

Exactly

OpenStudy (dumbsearch2):

your so smart! :)

OpenStudy (dumbsearch2):

thx for your help! :)

hero (hero):

You need to learn how to use a calculator bro.

hero (hero):

Or else you're going to get all your homework wrong.

OpenStudy (dumbsearch2):

Yeah I should have done the Order of Operations, I usually do that, don't know what's wrong with me! :| I usually do please excuse my dear aunt sally: http://www.webquest.hawaii.edu/kahihi/mathdictionary/images/PEMDAS.png :) well thanks for your help! :)

OpenStudy (dumbsearch2):

so like positive 7 right? :)

hero (hero):

Yes, bro. I already said it was correct.

hero (hero):

My only thing is, if you use a calculator, you have to know how to interpret what \(\large\frac{a \times b}{a + b}\) means

hero (hero):

Because when you see that, you must understand that it means to do the operations in the numerator and the denominator first before dividing.

hero (hero):

when you see that, dividing will be your final step.

OpenStudy (dumbsearch2):

right like in parenthesis right? like (axb)/(a+b) no?

hero (hero):

Exactly

hero (hero):

So \(\large\frac{a \times b}{a + b}\) implies that \(\large\frac{(a \times b)}{(a + b)}\)

hero (hero):

I'm sorry that your teacher neglected to tell you that.

OpenStudy (dumbsearch2):

ok. :)

OpenStudy (anonymous):

but the big division line bar also tells you that. or realizing that "a + b" is the denominator, it would never make sense to separate them in a computation. or, etc. etc. etc.

hero (hero):

I know @binarymimic, but I was just providing some meaningful clarification for the poor guy.

OpenStudy (anonymous):

you're so kind

OpenStudy (dumbsearch2):

thanks so much! :)